RC meeting and Run Plan, Thursday February 2nd

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Report


  • Swing shift, locked up Hall around 5pm.
  • Troubles with beam resotration. 2L25 cavity dropped HV during restore and then BLM trips ramping Hall A current during swing and owl.
  • Started moving on to beam checkout towards end of swing shift but had to stop ~4am due to the BLM issue.
  • Morning MCC meeting OPS say they will investigate BLM issues.
  • Day shift, we started harp scans ~11:30, started polhe3 running ~13:45

Current Status

  • Currently running polhe3, but beam not very stable.

Things to Discuss

  • SL had issues accessing site, despite all training completed. Something to check with users office. Sent reminder to upcoming shift crews and more generally halla_running to check access and training before coming for shift. Specifically GEN032 is causing issues.
  • Target oven, potentially tripped variac, may need opportunistic access to check.

Run Plan

  • Continue production running at 45uA and polHe3

Other Topics/Status Updates

  • Half wave plate. Hall B wants to change HWP, maybe tomorrow if get enough stats today. They will probably want to change every couple of days moving forward, depending on beam. I will find out more later.
  •  ?

Shift Schedule

Requested Accesses to the Hall

  • Looking to put SBS GEMs in DAQ during day shifts, GEM team need opportunistic access to hall if beam down
  • Target team for oven

Special Requests

  • DC power did not access Hall during maintenance day for moeller magnet, would be good to coordinate in the future.

Notices for Shift Crew

  • Check the How to page
  • Pay attention to the beam restoration procedures in the How To link
    • there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning. Communicate with MCC to figure out why beam has been down, to figure out which kind of restoration you should do.
  • SL:
    • production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
  • TOs:
    • NMR every 4 hours
    • TOs - please note the instructions for performing an NMR if there has been any changes to magnetic fields (NMR gui needs restarted, as well as changing the holding fields).
  • Contact RC any time for any questions

Data Collection and Polarisaton Status

  • Run List Sheet
  • Accumulated charge uncorrected 17.9C; accumulated charge corrected 18.654C (currently running HWP IN)
  • Accumulated charge uncorrected for OUT was 16.613C, accumulated charge corrected for OUT 17.433C)
  • Current target polarisation 52.2%

Useful Links